OpenStudy (anonymous):

Simplify the expression. -3/6-5/7

OpenStudy (anonymous):

find common denominator: find a least common product of 6 and 7

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

multiply cross wise: denominator down left with numerator up right, numerator up left with denominator down right

OpenStudy (anonymous):

can u just tell me the answer

OpenStudy (anonymous):

I dont know the answer yet this is how it goes: \[-\frac{3 }{ 6 } - \frac{ 5 }{ 7 }=-\frac{7 \times 3 }{ 7 \times 6 } - \frac{ 6 \times 5 }{ 6 \times 7 }\] do you understand what I did there?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

ok it is 21 - 30 over 42

OpenStudy (anonymous):

-21-30

OpenStudy (anonymous):

so -51 over 42 ?

OpenStudy (anonymous):

but my answer chooses say different

OpenStudy (anonymous):

this is the problem: \[-\frac{3}{6}-\frac{5}{7}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I get: -21/42 -30/42 6*7=42 3*7=21 5*6=30 I get -51/42. I don't know what wrong

OpenStudy (anonymous):

they wont it simplified

OpenStudy (anonymous):

want*

OpenStudy (anonymous):

you can still split away one whole: -1 9/42

OpenStudy (anonymous):

also: 10*3 = 30 and 4*3 = 12 so, 14*3 = 42 we get: 9/42 = 3 / 14 answer: -1 3/14